Let x represent the price of a suit: If the suit is on sale at a discount rate of 30%, express the price of the suit at the discount rate in terms of x.
Found 2 solutions by stanbon, nyc_function:
Answer by stanbon(75887)

Let x represent the price of a suit: If the suit is on sale at a discount rate of 30%, express the price of the suit at the discount rate in terms of x.
---------------------------------------
If you paid "x" for the suit, and you discounted it 30%,
the selling price would be:
S(x) = x - 0.30x = 0.7x
=============================
Cheers,
Stan H.

Answer by nyc_function(2741)

Let x = price of suit before the discount rate.
Let P = price of suit at a discount rate in terms of x.
This means P will EQUAL something on the other side of the equation where x is located. I used P but you can use any letters, of course.
P = x(0.30)